a circular oak table top is 4 ft. in diameter and 3/4 in. thick.How heavy is it if oak weighs 47 lb. per cu. ft.?
Answer by Theo(13342)

The oak table is 4 feet in diameter.
This means that the radius is 2 feet.
The oak table is 3/4 inches thick.
3/4 inches thick is equivalent to (3/4)/12 feet thick which equals .0625 feet thick.
The volume of the oak table is given by the formula

V = pi*r^2*h where:
pi = the constant of 3.14..... (it's in the calculator)
r = 2 feet
h = the thickness of the oak table = .0625 feet

The volume of the oak table is therefore

V = pi*2^2*.0625 = pi*4*.0625 = pi*.25 = .785398163 cubic feet.

If the oak table weighs 47 pounds per cubic feet, then the table top weighs 47 * .785398163 = 36.91371368 pounds.